OpenStudy (anonymous):

what is the simpliest form of the expression below? 3a/b x b^2/12C

OpenStudy (mitu12):

\[ab/4c\]

OpenStudy (kkutie7):

\[\frac{ 3a }{ b }*\frac{ b ^{2} }{ 12C }\] \[\frac{ 3ab ^{2}}{ 12bC }\] \[\frac{ ab }{ 4C }\]

OpenStudy (anonymous):

thanks

OpenStudy (kkutie7):

sure thing

OpenStudy (mitu12):

no problem